In an experiment the angles are required to be measured using an instrument. 29 divisions of the main scale exactly coincide with the 30 divisions of the vernier scale. If the smallest division of the main scale is half-a-degree (=0.5°), then the least count of the instrument is ?

Grade:11

2 Answers

Arun
3 years ago
We have, 
Least Count of Main Scale ( in degree) 
= 0.5° =1°/2 
2) According to the question, 
29 divisions of the main Scale coincide with 30 divisions on Vernier. 
=> 29 MSD = 30 VSD 
3) We know ,
Least Count, LC = 1 MSD - 1 VSD 
= 1 MSD - 29/30 MSD 
=1/30 MSD 
=1/30*1/2° =1°/60 
=1' = 1 minute 
Hence, Least Count of Instrument is 1 minute. 
 
Vikas TU
3 years ago
1) We have,
Least Count of Main Scale ( in degree)
= 0.5° =1°/2
2) According to the question,
29 divisions of the main Scale coincide with 30 divisions on Vernier.
=> 29 MSD = 30 VSD
3) We know ,
Least Count, LC = 1 MSD - 1 VSD
= 1 MSD - 29/30 MSD
=1/30 MSD
=1/30*1/2° =1°/60
=1' = 1 minute
Hence, Least Count of Instrument is 1 minute.